OpenStudy (mathmath333):

a test has 80 questions . there is one mark for a correct answer ,while there is negative penalty of -1/2 for a wrong answer and -1/4 for an unattempted question .what is the number of questions answered correctly, if the student has scored a net total of 34.5 marks

OpenStudy (anonymous):

If \(x,y,z\) are the numbers of questions answered correctly, incorrectly, or not answered (respectively), then you can set up the following system: \[\begin{cases}x+y+z=80\\ x-0.5y-0.25z=34.5 \end{cases}\] Are we missing any part to the question?

OpenStudy (mathmath333):

options a. 45 b. 48 c. 54 d. cannot be determined

OpenStudy (anonymous):

Ah alright then. You have two equations with three unknowns. The system has no particular solution; at best you can only solve for two variables in terms of the third variable.

OpenStudy (mathmath333):

cannot be determined?

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

but the book shows ans 54

OpenStudy (anonymous):

Either that's a typo, or the question is missing some info.

OpenStudy (mathmath333):

part b que. if the student has left 10 quetions unanswered ,the number of correct answers are a.45,b.48,c.54 d.cannot be determined

OpenStudy (anonymous):

10 unanswered questions means \(z=10\) in the equations above: \[\begin{cases}x+y+10=80\\ x-0.5y-0.25(10)=34.5 \end{cases}~~\iff~~\begin{cases}x+y=70\\ x-0.5y=32 \end{cases}\] This is a system with 2 equations and 2 unknowns. You can find a numerical value for \(x\).

OpenStudy (mathmath333):

48